# All Models Are Miscalibrated, But Some Less So: Comparing Calibration with Conditional Mean Operators

### v1

@@ 4.3 Joint Kernel Calibration Error
Marginal 0.001 6.908 $0$0.002 0.003
Figure 3: Distribution of calibration error estimates for four probabilistic models on CIFAR-10 and proportion of trials in which they are ranked correctly by calibration error.
5 Experimental Results
In our experiments, we assess the ability of the CKCE to distinguish between models through a relative comparison of calibration error. We compare its performance with ECE and JKCE. In all following experiments, the kernel bandwidth parameter $\gamma$ for both the CKCE and JKCE is chosen via the median heuristic. The ECE is estimated by uniform binning of the simplex, with each axis split into three equal intervals to ensure there are enough samples per bin, and we take $d$ to be the $L^{1}$ norm between probability vectors as is common in the literature Guo et al., (2017); Vaicenavicius et al., (2019).
5.1 Ranking Candidate Models
@@ 5.1.1 Synthetic Data
Synthetic data allows us to easily construct perfectly calibrated probabilistic predictive models. We consider a classification problem with $m=10$ classes and vary the number of samples $n$ used to measure calibration error. We construct the true conditional distribution by sampling from a Dirichlet distribution $P(Y|X=x_{i})\sim\text{Dir}(0.1,\dots,0.1)$, and then simulate true labels by $y_{i}\sim P(Y|X=x_{i})$ for $i=1,\dots,n$.
The two calibrated models we consider are the true conditional model, $f_{t}(x_{i})=P(Y|X=x_{i})$, and the marginal prediction model $f_{m}(x_{i})=(0.1,\dots,0.1)^{\top}$. To get miscalibrated models, we take the predictions of the true conditional distribution and either sharpen or soften the probability vector, producing an overconfident and underconfident model respectively. By defining miscalibrated predictive classifiers in such a way, we ensure that the top-class accuracy remains the same, but calibration error increases. The models in Figure 1 were constructed in a similar way, with more detail provided in Appendix C.1.
To assess the performance of the calibration error metrics, we do a pairwise comparison between each calibrated and miscalibrated model. This is repeated for 100 independent trials and we count the proportion of times that the estimated calibration error for the calibrated model is less than that of the micalibrated one for each metric. The results are displayed in Figure 2. As the number of samples in our calibration set increases, all three calibration error estimators are able to distinguish between the calibrated and miscalibrated models. However, when the number of calibration samples is small, CKCE generally has the most reliable performance.
(a) Synthetic Data
(b) Imagenet Data
5.1.2 Calibration of Neural Networks
We next consider ranking candidate models in a more practical setting. We used the CIFAR-10 dataset and three models with different neural network architectures: LeNet, DenseNet and ResNet. See Appendix C.3 for details on model hyperparameters and the training procedure. In addition, we also compared the marginal probability model which predicts equal probability across all labels.
Table 1 shows the accuracy, cross entropy and calibration error (CKCE, JKCE and ECE) for the models. CKCE and JKCE values are multiplied by 10 and 100 respectively so that the calibration errors are of comparable magnitude. We note that JKCE typically decreases in magnitude as the number of classes increases. The metrics were evaluated using the 10000 samples in the test set. ResNet shows the best performance in terms of accuracy and cross entropy, however the marginal model has the lowest calibration error as it is calibrated by definition. Using the full test set as a proxy for the ground truth calibration, the ranking of model calibration according to all three metrics is: 1. Marginal, 2. ResNet, 3. DenseNet, 4. LeNet.
To check the consistency of each metric, we randomly sample $n=500$ of the test set to form the calibration set and measure the calibration error for each of the models using the three estimators. We repeat for 1000 independent trials, and the resulting distribution of calibration error is presented via histograms in Figure 3. Additionally, we show the proportion of times each metric correctly ranks the models in the exact order as determined above.
The proportion indicates that out of the 1000 trials, the CKCE correctly ranked the four models by calibration 700 times. The histogram illustrates that the CKCE is the best metric at identifying the marginal probability model as the best calibrated. One advantage of the JKCE is that it has an unbiased estimator (which leads to the negative values of JKCE estimates, even though the population JKCE is a nonnegative quantity). On the other hand, CKCE and ECE estimators are both biased, which can be observed by comparing the distributions in Figure 3 with the values in Table 1. However, the variance of JKCE makes it unable to correctly rank the models nearly 50% of the time.
5.2 Covariate Shift
To demonstrate our claim that the CKCE is less sensitive to the marginal distribution of the predictive model, we now look at measuring calibration in problems with covariate shift. A toy example demonstrating the importance of this is presented in Appendix A.
5.2.1 Synthetic Data
@@
Set up: Let $\mathcal{X}\times\mathcal{Y}=[-1,1]\times\{0,1\}$ with a conditional distribution governed by $P(Y=1|X=x)=\frac{1}{1+e^{-x}}$. Next, we choose a (miscalibrated) probabilistic model $Q$ with rule $Q(Y=1|X=x)=\frac{1}{1+e^{-5x}}$. We let the marginal distribution $P(X)$ be a truncated normal with domain $[-1,1]$, scale parameter $\sigma=0.25$ and location parameter $\alpha$ which we vary between $-1$ and $1$. Note that the conditional distribution $P(Y|X)$ remains unchanged, making this an example of covariate shift. The set up is illustrated in Figure 4(a) (left). At each value of $\alpha$, we sample $n=1000$ inputs from $P(X)$ and simulate labels using $P(Y|X)$. The resulting calibration set is used to estimate each of CKCE, JKCE and ECE. Unlike in Section 5.1.2, these values are not scaled.
What we expect: Even though a change in $P(X)$ changes the marginal distribution of $Q_{X}$, the conditional distribution $P(Y|Q_{X})$ and $P(Z_{X}|Q_{X})$ are not effected. Therefore, we would expect that calibration error stays constant as the location parameter $\alpha$ changes. This would indicate that the calibration error estimator is robust against covariate shift.
@@
What we observe: The results in Figure 4(a) (right) show that the CKCE estimator indeed stays constant, with some scattering due to randomness in the generated samples, demonstrating the desired behaviour. On the other hand, both JKCE and ECE are highly sensitive to changes in $P(X)$, showing clear variability as the parameter $\alpha$ shifts.
@@ 5.2.2 ImageNet
Set up: We next evaluate the impact of distribution shift on the calibration of image models, using ResNet and ViT neural networks trained on ImageNet. See Appendix C.3 for model details. Their performance on the full validation set is displayed in Table 1. We take a subset of $n=1000$ validation data, and alter the brightness of the input image, ranging from -50% to +50% in increments of 10%. At each brightness level, the CKCE and JKCE for both models is computed. Due to the large number of classes, ECE binning scheme undergoes a combinatorial explosion and cannot be accurately estimated.
What we expect: As the brightness changes, we would expect that model miscalibration, and hence relative ranking of the models, remains fairly constant. While an alteration in image brightness may cause a shift in the marginal distribution of the predictive models, the conditional distribution $P(Y|Q_{X})$ should be stable.
What we observe: Figure 4(b) reveals that CKCE consistently indicates that ViT has lower calibration error than ResNet, with the estimate for both models largely unaffected by the change in input image brightness. This agrees with the findings of Minderer et al., (2021) which showed that transformer based models typically have better calibration than those with convolutional layers. On the other hand, JKCE is highly sensitive to the shift in input distribution, with the estimator preferring different models for different brightnesses.
@@ 6 Conclusion
We introduced the CKCE - a novel method of measuring calibration based on conditional mean operators. We demonstrated that the CKCE provides a more reliable way of performing a relative comparison of calibration of probabilistic predictive models compared to other metrics. This allows it to better distinguish calibrated models from miscalibrated ones, and more accurately rank a collection of models. Furthermore, it remains stable under distribution shift. While we do not address ways to improve model calibration in this work, future research could explore using CKCE as a regularization term in model training. Its ability to more precisely capture relative comparisons as training progresses may offer advantages over other metrics.
Appendix
Appendix A Effect of marginal distribution $P(Q_{X})$
A change in the marginal distribution of the predictive model will cause a change in the joint distribution $P(Y,Q_{X})$, even if the conditional $P(Y|Q_{X})$ remains the same. That is, $P(Y,Q_{X})=P(Y|Q_{X})P(Q_{X})\neq P(Y|Q_{X})P^{\prime}(Q_{X})=P^{\prime}(Y,Q_{X})$.
The effect this has on the calibration of a model can be illustrated with the following toy example. Consider a setting with $\mathcal{X}\times\mathcal{Y}=\{0,1\}\times\{0,1\}$ and a simple probabilistic classifier $Q$ as summarised in Table 2. Clearly, the model is miscalibrated, however the extent of this depends on which calibration measure is chosen.
Table 2: Summary of a toy model
$x$ $P(X=x)$ $Q(Y=1|X=x)$ $P(Y=1|X=x)$
0 $p$ 0.3 0.5
1 $1-p$ 0.6 0.5
To calculate CKCE, we will choose kernels $\ell(y,y^{\prime})=\mathbf{1}\{y=y^{\prime}\}$ and $k(q,q^{\prime})=\mathbf{1}\{q=q^{\prime}\}$. Since the variables are discrete, CMOs become conditional probability matrices, allowing them to be written explicitly Song et al., (2013), e.g.
$\mathcal{C}_{Y|Q_{X}}=[P(Y=y|Q_{X}=q)]_{y\in\{0,1\},q\in\{0.3,0.6\}}.$
Furthermore, the Hilbert-Schmidt norm is simply the Frobenius norm, meaning CKCE is easily computed.
$\text{CKCE}_{k,\ell}=\|\mathcal{C}_{Y|Q_{X}}-\mathcal{C}_{Z_{X}|Q_{X}}\|_{HS}^{2}=\left\|\begin{bmatrix}0.5&0.5\\ 0.5&0.5\end{bmatrix}-\begin{bmatrix}0.7&0.4\\ 0.3&0.6\end{bmatrix}\right\|_{F}^{2}=0.1.$
Using the same kernels as above to define $\kappa=\ell\otimes k$, we will compute the JKCE. Now the joint embeddings can also be represented explicitly, this time as joint probability matrices, e.g.
$\mu_{YQ_{X}}=[P(Y=y\cap Q_{X}=q)]_{y\in\{0,1\},q\in\{0.3,0.6\}}.$
Again, the RKHS norm is the Frobenius norm, and so
$\text{JKCE}_{\kappa}=\|\mu_{YQ_{X}}-\mu_{Z_{X}Q_{X}}\|_{HS}^{2}=\left\|\begin{bmatrix}0.5p&0.5(1-p)\\ 0.5p&0.5(1-p)\end{bmatrix}-\begin{bmatrix}0.7p&0.4(1-p)\\ 0.3p&0.6(1-p)\end{bmatrix}\right\|_{F}^{2}=0.1p^{2}-0.04p+0.02.$
Unlike for CKCE, which was not impacted by the marginal distribution of $X$, JKCE depends on the parameter $p$. A similar problem occurs with ECE. Choosing metric $d(x,y)=\|x-y\|_{L_{1}}$, gives
$\text{ECE}_{d}$ $=\mathbb{E}_{X}d(P(Y|Q_{X}),Q_{X})$
$=(1-p)d((0.5,0.5)^{\top},(0.7,0.3)^{\top})+pd((0.5,0.5)^{\top},(0.4,0.6)^{\top})$
$=0.4-0.2p$
As with JKCE, the level of miscalibration according to ECE depends on the parameter $p$. Having the calibration of a model be dependent on the marginal distribution of the inputs can be harmful in certain scenarios. For example, let $X$ indicate whether a person is a man or a woman, and $Y$ indicate whether they are a good hire for a company. Thus, evaluating a model’s calibration using JKCE or ECE would be heavily impacted by who the company choose to interview (which determines the calibration set), causing fairness concerns. On the other hand, this is not an issue for CKCE, which only compares the conditional distributions.
Appendix B Random Fourier Feature Implementation
In Section 3.2, we propose the kernel $k:\Delta^{d-1}\times\Delta^{d-1}\to\mathbb{R}$ defined by
$k(p,p^{\prime})=p^{\top}p^{\prime}+\exp\left(-\frac{1}{2\gamma^{2}}\|p-p^{\prime}\|^{2}\right),$
along with $\ell(y,y^{\prime})=\mathbf{1}\{y=y^{\prime}\}$ on the space of labels for measuring calibration of probabilistic classifiers using the CKCE. We comment on how random Fourier features can be used to approximate the feature mapping for $k$ allowing us to compute CMO estimators in the primal. Here we provide further detail for practical implementation of this method.
Note that adding kernels is equivalent to concatenating features. Thus, an approximation of the canonical feature mapping for kernel $k$ is
$\hat{\phi}(p)=\left[p^{(1)},\dots,p^{(d)},\frac{1}{\sqrt{D}}\cos(w_{1}^{\top}p),\frac{1}{\sqrt{D}}\sin(w_{1}^{\top}p),\dots,\frac{1}{\sqrt{D}}\cos(w_{D}^{\top}p),\frac{1}{\sqrt{D}}\sin(w_{D}^{\top}p)\right]^{\top}$
where $p=[p^{(1)},\dots,p^{(d)}]^{\top}\in\Delta^{d-1}$ and $w_{1},\dots,w_{D}$ are $D$ independent samples from the multivariate normal distribution $\mathcal{N}_{d}(0,\gamma^{-2}I)$. One can show that $\mathbb{E}[\hat{\phi}(p)^{\top}\hat{\phi}(q)]=k(p,q)$ and $\text{Var}[\hat{\phi}(p)^{\top}\hat{\phi}(q)]=O(D^{-1})$ (Rahimi and Recht,, 2007). In our experiments, we choose $D=100$.
Using the approximation $\hat{\phi}$ rather than the full infinite dimensional feature mapping allows to estimate each CMO without calculating a Gram matrix. Given a calibration set $\{(y_{i},q_{i})\}_{i=1}^{n}$, where each $q_{i}$ is the probability vector predicted by our classifier for a feature $x_{i}$, the CMO estimators are
$\widehat{\mathcal{C}}_{Y|Q_{X}}$ $=\Psi_{Y}\widehat{\Phi}_{Q}^{*}(\widehat{\Phi}_{Q}\widehat{\Phi}_{Q}^{*}+\lambda nI_{m})^{-1}$
$\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}$ $=M_{Q}\widehat{\Phi}_{Q}^{*}(\widehat{\Phi}_{Q}\widehat{\Phi}_{Q}^{*}+\lambda nI_{m})^{-1}$
where $\Psi_{Y}=[\psi(y_{1})\dots\psi(y_{n})]$, $M_{Q}=[q_{1},\dots q_{n}]$, $\widehat{\Phi}_{Q}=[\hat{\phi}(q_{1})\dots\hat{\phi}(q_{n})]$, $\lambda$ is a regularisation parameter and $I_{m}$ is the $m\times m$ identity matrix with $m=d+2D$. The resulting CKCE estimator is
$\widehat{\text{CKCE}}_{k,\ell}=||\widehat{\mathcal{C}}_{Z_{X}|Q_{X}}-\widehat{\mathcal{C}}_{Y|Q_{X}}||_{F}^{2}$
where we have replaced the Hilbert-Schmidt norm with the Frobenius norm as we are now working with finite dimensional matrices rather than operators. As the main computational complexity comes from calculating the inverse, we have thus reduced the computational cost from $O(n^{3})$ to $O(m^{3})$ with $m\ll n$.
The theoretical analysis of random Fourier features has been well studied (cf. Li et al., (2021) and references therein), demonstrating guaranteed computational savings without sacrificing convergence properties of estimators in many cases.
Appendix C Experimental Details
C.1 Synthetic Miscalibration
In Section 5.1.1, we perform a relative comparison of a calibrated and miscalibrated model. Constructing well calibrated models from the synthetic data is relatively straightforward, simply taking the true conditional and marginal probability distributions.
  • True Conditional: $f_{t}(x_{i})=P(Y|X=x_{i})$,
  • Marginal: $f_{m}(x_{i})=(0.1,\dots,0.1)^{\top}$.
To produce miscalibrated models, we take the true conditional distribution and either sharpen or soften the probability vector by temperature scaling by a parameter $t\in[0,\infty)$. Given a probability vector $p\in\Delta^{m-1}$, we get its scaled version by $p_{t}=\text{softmax}(\text{clr}(p)/t)$, where $\text{clr}:\Delta^{m-1}\to\mathbb{R}$ is the centre log ratio transformation (Aitchison,, 1982).
Since softmax and clr are inverses of each other, setting the temperature to $t=1$ returns the probability vector unchanged. Having $t>1$ softens $p$, whereas $t<0$ sharpens $p$. Thus, the two miscalibrated model are
  • Underconfident: $f_{u}(x_{i})=\text{softmax}(\text{clr}(f_{t}(x_{i}))/2)$,
  • Overconfident: $f_{o}(x_{i})=\text{softmax}(\text{clr}(f_{t}(x_{i}))/0.5)$.
C.2 Reliability Diagram
Figure 1 in Section 1 shows a reliability diagram for three probabilistic predictive models. Here we provide extra detail on how the models and resulting diagram are generated. The models are binary probabilistic classifiers, that is $m=2$, and $n=1000$ synthetic samples were generated for the calibration set. The synthetic data is generated by first sampling true conditional distributions from a Dirichlet distribution $P(Y|X=x_{i})\sim\text{Dir}(1,1)$, and then simulating true labels by $y_{i}\sim P(Y|X=x_{i})$ for $i=1,\dots,n$, similar to the set up in Section 5.1.1.
As in Section 5.1.1, the true conditional model is simply
True Conditional: $\text{{True Conditional:} }f_{t}(x_{i})=P(Y|X=x_{i})$
which is perfectly calibrated by definition. The two miscalibrated models are
Underconfident: $\text{{Underconfident:} }f_{u}(x_{i})=\text{softmax}(\text{clr}(f_{t}(x_{i})/2))$
Overconfident: $\text{{Overconfident:} }f_{o}(x_{i})=\text{softmax}(\text{clr}(f_{t}(x_{i})/0.5))$
where we have used temperature scaling to either soften or sharpen the true probability vectors as described above.
The calibration set is formed by the pairs $\{(y_{i},f(x_{i}))\}_{i=1}^{n}$ for each of the models. To produce the reliability diagram, we first partition the interval $[0,1]$ into 10 uniform bins with $B_{m}$ containing the set of indices for which $f^{(0)}(x_{i})\in[\frac{m-1}{10},\frac{m}{10}]$ for $m=1,\dots 10$, that is, the predicted confidence in class 0 falls in bin $m$. Then, the average confidence in bin $m$ is given by
$\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{k\in B_{m}}f^{(0)}(x_{k}),$
while the average accuracy is
$\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{k\in B_{m}}\mathbf{1}\{y_{k}=0\}.$
Plotting the confidence and accuracy value in each bin for each model produces the reliability diagram. The CKCE value in Figure 1 is computed using the same 1000 samples used to produce the reliability diagram.
C.3 Neural Network Architecture
In Section 5.1.1 we conduct experiments using three pretrained neural networks obtained from Kull et al., (2019) on the CIFAR-10 dataset [^2] Source code and output logits available at https://github.com/markus93/NN_calibration. Below we provide further details for each model.
LeNet CIFAR-10
We used a standard LeNet-5 convolutional neural network architecture (Lecun et al.,, 1998). The model is trained for 300 epochs using stochastic gradient descent with cross entropy loss.
DenseNet CIFAR-10
We used a standard DenseNet-40 convolutional neural network architecture (Huang et al.,, 2018). The model is trained for 300 epochs using stochastic gradient descent with cross entropy loss.
ResNet CIFAR-10
We used a standard ResNet-110 convolutional neural network architecture (He et al.,, 2015). The model is trained for 200 epochs using stochastic gradient descent with cross entropy loss.
In Section 5.2.2 we used the ImageNet dataset, and pretrained models available in the timm library.
ResNet ImageNet
We used a ResNet-50 convolutional neural network architecture (He et al.,, 2015) with the resnet50.a1_in1k parameters from the timm library.
Vit ImageNet
We used a Vision Transformer neural network architecture (Zhai et al.,, 2023) with the ViT-L-16-SigLIP-384 parameters from the timm library.

### v2

@@ 5 Experimental Results
Marginal 0.001 6.908 0.000 -0.002 0.003
5.1 Robustness Against Differences in Marginal Distribution of Predictions
We begin by comparing the calibration error of models with different marginal distribution of predictions on the ImageNet dataset. We consider two pretrained neural network image classifiers, ResNet and ViT available from the timm library [^2] The specific parameters used are resnet50.a1_in1k and ViT-L-16-SigLIP-384. In addition, we also compared the marginal probability model which predicts equal probability across all labels. Table 1 shows the accuracy, cross entropy and calibration error (CKCE, JKCE and ECE) for the models computed on the validation data. Although all calibration error metrics indicate that the marginal distribution model has the lowest calibration error (JKCE produces a negative value as it is an unbiased estimator), there is a disagreement regarding which neural network is best calibrated. CKCE indicates that the ViT model has lower calibration error which agrees with the findings of [Minderer_2021_revisiting] that transformer based models typically have better calibration than those with convolutional layers. On the other hand, JKCE ranks the models equally while ECE has ResNet as superior.
Figure 1 provides more insight into why this occurs. The histograms show that most ResNet predictions lie in the corners of the probability simplex, whereas ViT is more conservative. Indeed the reliability diagrams in the same plot show that ResNet is overconfident, while ViT is slightly underconfident. Despite the reliability diagram for ViT appearing to have a closer match between model confidence and accuracy, particularly in the interval [0.5, 1], ResNet is preferred in terms of ECE as it makes more predictions in regions of better calibration. It is in some sense rewarded for its overconfidence. However, this may not be ideal for practitioners who would like for all outputs to be well calibrated. In contrast, our CKCE metric is unaffected by the marginal distribution of outputs, and thus provides a more reliable measure of calibration.
5.2 Robustness Against Covariate Shift
@@ 5.2.1 Synthetic Data
Let $\mathcal{X}\times\mathcal{Y}=[-1,1]\times\{0,1\}$ with a conditional distribution governed by $P(Y=1|X=x)=\frac{1}{1+e^{-x}}$. Next, we choose a (miscalibrated) probabilistic model $f(x)=\frac{1}{1+e^{-5x}}$. We let the marginal distribution $P(X)$ be a truncated normal with domain [1,1][-1,1], scale parameter $\sigma=0.25$ and location parameter $\alpha$ which we vary between 1-1 and $1$. Note that the conditional distribution $P(Y|X)$ remains unchanged, making this an example of covariate shift. The set up is illustrated in Figure 2 (left). At each value of $\alpha$, we sample $n=1000$ inputs from $P(X)$ and simulate labels using $P(Y|X)$. The resulting calibration set is used to estimate each of CKCE, JKCE and ECE.
@@
Even though a change in $P(X)$ changes the marginal distribution of $Q_{X}$, the conditional distribution $P(Y|Q_{X})$ and $P(Z_{X}|Q_{X})$ are not effected. Therefore, we would like for that calibration error to stay constant as the location parameter $\alpha$ changes. This would indicate that the calibration error estimator is robust against covariate shift. The results in Figure 2 (right) show that the CKCE estimator indeed stays constant, with some scattering due to randomness in the generated samples, demonstrating the desired behaviour. On the other hand, both JKCE and ECE are highly sensitive to changes in $P(X)$, showing clear variability as the parameter $\alpha$ shifts.
@@ 5.2.2 ImageNet
We next evaluate the impact of distribution shift on the calibration of image models, using the ResNet and ViT from Section 5.1. Their performance on the full validation set is displayed in Table 1. Here we take a subset of $n=1000$ validation data, and alter the brightness of the input image, ranging from -50% to +50% in increments of 10%. At each brightness level, the CKCE and JKCE for both models is computed. Due to the large number of classes, the ECE binning scheme suffers from the curse of dimensionality and cannot be accurately estimated.
As the brightness changes, we would expect that model miscalibration, and hence relative ranking of the models, remains fairly constant. While an alteration in image brightness may cause a shift in the marginal distribution of $Q_{X}$, the conditional distribution $P(Y|Q_{X})$ should be stable. Indeed, Figure 3 shows that CKCE consistently indicates that ViT has lower calibration error than ResNet, with the estimate for both models largely unaffected by the change in input image brightness. On the other hand, JKCE is highly sensitive to the shift in input distribution, with the estimator preferring different models for different brightness levels.
Figure 3: Calibration error of models with changing image brightness. CKCE provides consistent model preference, unlike JKCE.
5.3 Kernel choice robustness
We conclude with an experiment that demonstrates the necessity for the choice of kernel in Section 3.2. Using the same synthetic data setting as in Section 5.2.1, we compare the kernel (8) against simply using a linear or Gaussian kernel. As before, we range the parameter $\alpha$ between -1 and 1 at 25 discrete values and calculate the CKCE using the three choices of kernel $k$. The mean of 20 trials is plotted in Figure 4.
As we saw in the previous section, computing the CKCE with the combined linear-Gaussian kernel provides a metric that is robust against covariate shift. On the other hand, when a linear or Gaussian kernel are used individually on the input variable, then the CKCE no longer remains consistent, with a noticeable drop near $\alpha=0$. The calibration error plot suggests that kernel (8) leads to CMO estimators that are less sensitive to the marginal distribution of inputs.
@@ 6 Conclusion
We introduced the CKCE - a novel method of measuring calibration based on conditional mean operators. We demonstrated that the CKCE provides a more reliable way of performing a relative comparison of calibration of probabilistic classifiers compared to other metrics. It is less effected by differences in marginal distributions of model outputs and remains stable under covariate shift. While we do not address ways to improve model calibration in this work, future research could explore using CKCE as a regularisation term in model training. Its ability to more precisely capture relative comparisons as training progresses may offer advantages over other metrics.
Figure 4: CKCE of a probabilistic model for three choices of kernel on the input variable. Using a kernel that combines a linear and a Gaussian component provides a more robust measure of calibration error.
{credits}
6.0.1 Acknowledgements
The authors would like to express their gratitude to Jiawei Liu and Russell Tsuchida for providing helpful guidance and discussion in the early stages of this work.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.